A gas turbine operating on Brayton cycle has the maximum temperature of 1200K and the
minimum temperature of 300K. The cycle efficiency for the maximum work capacity
will be
 

Grade:12th pass

3 Answers

Vikas TU
6 years ago
Maximum Temperature = 1200 K
Minimum Temperature = 300K
Then , maximum efficiency = 1- Tmin/Tmax
=  1- 300/1200
= 1- ¼
= ¾ = 75%
Maximum eff = 75% minimum eff =25%
Taaggg
4 years ago
50% hoga.. kaise ye made easy walo se puchho jinhone is question ka answer dena dhi ni smjha.. chutiya made easy.. 
 
Prince Chaudhary
2 years ago
Cycle efficiency for the maximum work capacity of a gas turbine operating in Bratton cycle is;
 
     =  1-( min. Temp/ max. Temp.)^0.5
      =1-(300/1200)^0.5
      =0.50 or 50%